# HTMA-Net: Towards Multiplication-Avoiding Neural Networks via Hadamard Transform and In-Memory Computing

### v1

@@ Abstract
Reducing the cost of multiplications is critical for efficient deep neural network deployment, especially in energy-constrained edge devices. In this work, we introduce HTMA-Net, a novel framework that integrates the Hadamard Transform (HT) with multiplication-avoiding (MA) SRAM-based in-memory computing to reduce arithmetic complexity while maintaining accuracy. Unlike prior methods that only target multiplications in convolutional layers or focus solely on in-memory acceleration, HTMA-Net selectively replaces intermediate convolutions with Hybrid Hadamard-based transform layers whose internal convolutions are implemented via multiplication-avoiding in-memory operations. We evaluate HTMA-Net on ResNet-18 using CIFAR-10, CIFAR-100, and Tiny ImageNet, and provide a detailed comparison against regular, MF-only, and HT-only variants. Results show that HTMA-Net eliminates up to 52% of multiplications compared to baseline ResNet-18, ResNet-20, and ResNet-50 models, while achieving comparable accuracy in evaluation and significantly reducing computational complexity and the number of parameters. Our results demonstrate that combining structured Hadamard transform layers with SRAM-based in-memory computing multiplication-avoiding operators is a promising path towards efficient deep learning architectures.
@@ 1 Introduction
The rapid growth of deep neural networks (DNNs) has intensified the demand for energy-efficient and high-throughput hardware acceleration [1], [2]. In particular, multiply accumulate (MAC) operations dominate the workload in convolutional neural networks (CNNs), leading to significant energy consumption [3], [4], [5],[6]. Since multiplication is orders of magnitude more costly than addition in digital hardware [7], reducing multiplications is a key strategy for energy-efficient DNN inference. Traditional digital processors struggle to maintain the memory and computation requirements of modern inference workloads, especially in resource-constrained deployment scenarios. This challenge is amplified by the well-known memory wall, where the cost of data movement far exceeds that of arithmetic operations, creating a fundamental bottleneck in both latency and energy consumption.
@@ 2.1 Hadamard Transform
The Walsh-Hadamard Transform (WHT) is an orthogonal linear transform widely used in signal processing and, more recently, in efficient neural network architectures. Let $H_{m}\in\mathbb{R}^{2^{m}\times 2^{m}}$ denote the Hadamard matrix of order $2^{m}$, defined recursively as
@@ 4.1 Experimental Setup
We evaluate HTMA-Net on three benchmark datasets: CIFAR-10, CIFAR-100, and a subset of Tiny-ImageNet. All models are trained from scratch using standard SGD with momentum, cosine learning rate decay, and standard data augmentation (random cropping, horizontal flipping). ResNet-18 serves as the primary baseline, with ResNet-20 and ResNet-50 included to study scalability.
@@ 4.2 Experimental Results
Table 2 extends the analysis to different input resolutions and deeper networks (ResNet-20 and ResNet-50). We compare the baseline, MF-only, HT-only, and hybrid HTMA variants under the all-stages integration. Although MF-only yields the largest multiplication reduction, it suffers severe accuracy degradation as Fig. 1). By contrast, HTMA achieves $\geq 51\%$ multiplication elimination with negligible accuracy loss, consistently outperforming MF-only and matching or exceeding HT-only.

### v2

@@ Abstract
Reducing the cost of multiplications is critical for efficient deep neural network deployment, especially in energy-constrained edge devices. In this work, we introduce HTMA-Net, a novel framework that integrates the Hadamard Transform (HT) with multiplication-avoiding (MA) SRAM-based in-memory computing to reduce arithmetic complexity while maintaining accuracy. Unlike prior methods that only target multiplications in convolutional layers or focus solely on in-memory acceleration, HTMA-Net selectively replaces intermediate convolutions with Hybrid Hadamard-based transform layers whose internal convolutions are implemented via multiplication-avoiding in-memory operations. We evaluate HTMA-Net on ResNet-18 using CIFAR-10, CIFAR-100, and Tiny ImageNet, and provide a detailed comparison against regular, MF-only, and HT-only variants. Results show that HTMA-Net eliminates up to 52% of multiplications compared to baseline ResNet-18, ResNet-20, and ResNet-32 models, while achieving comparable accuracy in evaluation and significantly reducing computational complexity and the number of parameters. Our results demonstrate that combining structured Hadamard transform layers with SRAM-based in-memory computing multiplication-avoiding operators is a promising path towards efficient deep learning architectures.
@@ 1 Introduction
The rapid growth of deep neural networks (DNNs) has intensified the demand for energy-efficient and high-throughput hardware acceleration [1, 2]. In particular, multiply accumulate (MAC) operations dominate the workload in convolutional neural networks (CNNs), leading to significant energy consumption [3, 4, 5, 6]. Since multiplication is orders of magnitude more costly than addition in digital hardware [7], reducing multiplications is a key strategy for energy-efficient DNN inference. Traditional digital processors struggle to maintain the memory and computation requirements of modern inference workloads, especially in resource-constrained deployment scenarios. This challenge is amplified by the well-known memory wall, where the cost of data movement far exceeds that of arithmetic operations, creating a fundamental bottleneck in both latency and energy consumption.
@@ 2.1 Hadamard Transform
The Walsh-Hadamard Transform (WHT) is an orthogonal linear transform widely used in signal processing and, more recently, in efficient neural network architectures. As shown in Fig.2(a), the WHT folow a similar structure as Fast Fourier transform (FFT) and Wavelets [16, 17, 18, 19].
(a) WHT
(b) Reduced WHT
Fig. 2: a) Flow diagram of the N=8 fast Walsh-Hadamard transform, b) upper diagonal of the Walsh-Hadamard flow diagram.
Let $H_{m}\in\mathbb{R}^{2^{m}\times 2^{m}}$ denote the Hadamard matrix of order $2^{m}$, defined recursively as
@@ 3.1 Hadamard–Transform Layer
As a direction for future work, we plan to explore reduced or partial WHT variants, as illustrated in Fig. 2(b), which may further decrease arithmetic complexity within the transform stage while maintaining performance.
@@ 4.1 Experimental Setup
We evaluate HTMA-Net on three benchmark datasets: CIFAR-10, CIFAR-100, and a subset of Tiny-ImageNet. All models are trained from scratch using standard SGD with momentum, cosine learning rate decay, and standard data augmentation (random cropping, horizontal flipping). ResNet-18 serves as the primary baseline, with ResNet-20 and ResNet-32 included to study scalability.
@@ 4.2 Experimental Results
Table 2 extends the analysis to different input resolutions and deeper networks (ResNet-20 and ResNet-32). We compare the baseline, MF-only, HT-only, and hybrid HTMA variants under the Middle-only stage integration. Although MF-only yields the largest multiplication reduction, it suffers severe accuracy degradation as Fig. 1). By contrast, HTMA achieves $\geq 51\%$ multiplication elimination with negligible accuracy loss, consistently outperforming MF-only and matching or exceeding HT-only.